Solve the system by addition.
5x - 3y = 13
4x - 3y = 11

Answer by checkley77(12844)

5x - 3y = 13
-(4x - 3y = 11)
-----------------------
X=2 ANSWER.
5*2-3Y=13
10-3Y=13
-3Y=13-10
-3Y=3
Y=3/-3
Y=-1 ANSWER.